# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502 Japan

### v1

@@ 1 Introduction
Our main finding is that the measured critical exponents of the dipolar fixed points can be very different from those of the Heisenberg fixed points: $\nu=0.601(2)(^{+0}_{-2})$ and $\eta=0.132(8)(^{+0}_{-4})$ at $\lambda=8$, which should be compared with $\nu=0.712(1)(^{+3}_{-0})$ and $\eta=0.0318(27)(^{+10}_{-7})$ at $\lambda=0$. In Figure 1, we compare our results of $\nu$ with the previous Monte Carlo studies Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz as well as the conformal bootstrap Chester:2020iyt and the $\varepsilon$-expansion 2022NuPhB.98515990K . As for $\lambda=0$, our result is almost consistent with the other numerical works, while for $\lambda=8$ it shows a clear difference from the one at $\lambda=0$. Thus, it strongly suggests that the dipolar fixed point realized at $\lambda=\infty$ would be in a different universality class from the Heisenberg fixed point.
Figure 1: Comparison plot on the critical exponent $\nu$ (for $\lambda=0$ and $8$) obtained in this paper and the previous studies by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz (colored open symbols), the conformal bootstrap Chester:2020iyt (star-yellow), and the $\varepsilon$-expansion (filled symbols). The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion obtained for $\lambda=0$ and $\infty$ in Ref. Aharony_1 and for $\lambda=8$ in this paper whereas the filled square-circle ones are the 3-loop results taken from Ref. 2022NuPhB.98515990K .
@@ 4.1 Effectiveness of the local dipolar constraint
We measure the expectation values of the Heisenberg part, $A_{n}:=\sum_{i}\expectationvalue*{\vec{S}_{n}\cdot\vec{S}_{n+\hat{i}}}$, and the local dipolar constraint term, $B_{n}:=\expectationvalue*{(\vec{\nabla}_{n}\cdot\vec{S}_{n})^{2}}$, of the Hamiltonian (14), using generated Monte Carlo samples. Here, we take the lattice size $L=40$ and set the temperature to $\beta=0.69019$, $0.35288$, and $0.320374$ for $\lambda=0$, $4$, and $8$, respectively. These values of $\beta$ are indeed the peak positions of the magnetic susceptibility $\chi_{m}$. (See Appendix C for the details of these choices.)
@@
The average values of them over the whole lattice sites are shown in Table 1. We find that the local dipolar constraint term and the Heisenberg term give the same order of contribution to the Hamiltonian given by $H=\sum_{n}(-A_{n}+\lambda B_{n})$. This means that larger $\lambda$ gives the more strict imposition of the constraint as expected. Furthermore, the results for $\lambda=4$ and $8$ are roughly consistent with the virial theorem.
We further investigate a local property of the dipolar constraint. In the Hamiltonian (14), the dipolar constraint term forces the spin $\vec{S}_{n}$ to satisfy $(\vec{\nabla}_{n}\cdot\vec{S}_{n})^{2}=0$ at each site $n$. Here, we measure $B_{n}$ locally at each site $n=(x,0,0)$ as depicted in Figure 2 for $\lambda=0$ (circle-blue symbol), $4$ (square-green), and $8$ (diamond-red).
Figure 2: The results of $B_{n}=\expectationvalue*{(\vec{\nabla}_{n}\cdot\vec{S}_{n})^{2}}$ are plotted against the coordinate $x$ on the $x$-axis $n=(x,0,0)$. The lattice size is set to $L=40$. The symbols of circle-blue, square-green, and diamond-red correspond to $\lambda=0$, $4$, and $8$, respectively. The colored horizontal lines depict the average values over the whole lattice sites.
@@ 5.4 Critical exponent: $\nu$
These results of $\nu$ above are compared with the results of previous studies with various methods as shown in Figure 1. Here, we compare the values of $\nu$ for $\lambda=0$, $8$, and $\infty$ obtained by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz , the conformal bootstrap Chester:2020iyt , and the $\varepsilon$-expansion Refs. Aharony_1 ; 2022NuPhB.98515990K . For reference, the conformal bootstrap predicts $\nu=0.7117(2)$ 444 Here, the error in the bootstrap bound is not one-sigma range but rigorous Chester:2020iyt . and the latest Monte Carlo study Hasenbusch:2020pwj reports $\nu=0.71164(10)$ for the Heisenberg model. Our result at $\lambda=0$ is consistent with them. On the other hand, the result at $\lambda=8$ shows a large deviation from the analytical prediction by the 1-loop $\varepsilon$-expansion. Here, we note that the difference between the 1-loop result of the $\varepsilon$-expansion and the 3-loop result at $\lambda=0$ is larger than the one between the 1-loop one at $\lambda=0$ and $\lambda=8$. It indicates that loop corrections are sizable, at least larger than the $\lambda$-dependence.
@@ 7 Summary and discussion
Finally, we wish there would be a continuous effort to experimentally verify the dipolar effects in real ferromagnets. [^6] In the Chinese character, the root of magnetism is associated with a stone with benevolence. Why? In ancient Chinese stories, it is said that the way the ferromagnet attracts iron rocks makes it look as if an infant is naturally attracted to his or her mother. Probably the dipolar magnet will embody more benevolence because the magnetism from it is divergenceless, so it appears to be more gentle and embracing. So far in human history, the distinct features of dipolar magnetism have been observed in esoteric materials such as Europium compounds, but the universal nature of the benevolence should be also seen in more conventional materials such as iron or nickel in future experiments, just with more care.
@@ Acknowledgements.
The numerical calculations were carried out on Yukawa-21 at YITP in Kyoto University and the PC clusters at RIKEN iTHEMS. The work of E. I. is supported by JST PRESTO Grant Number JPMJPR2113, JSPS Grant-in-Aid for Transformative Research Areas (A) JP21H05190, JST Grant Number JPMJPF2221, JPMJCR24I3, and also supported by Program for Promoting Researches on the Supercomputer “Fugaku” (Simulation for basic science: from fundamental laws of particles to creation of nuclei) and (Simulation for basic science: approaching the new quantum era), and Joint Institute for Computational Fundamental Science (JICFuS), Grant Number JPMXP1020230411. The work by YN is in part supported by JSPS KAKENHI Grant Number 21K03581. This work is supported by Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics.

### v2

@@ 1 Introduction
Our main finding is that the measured critical exponents of the dipolar fixed points can be very different from those of the Heisenberg fixed points: $\nu=0.601(2)(^{+0}_{-2})$ and $\eta=0.132(8)(^{+0}_{-4})$ at $\lambda=8$, which should be compared with $\nu=0.712(1)(^{+3}_{-0})$ and $\eta=0.0318(27)(^{+10}_{-7})$ at $\lambda=0$. In Figure 1, we compare our results of $\nu$ with the previous Monte Carlo studies Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz as well as the conformal bootstrap Chester:2020iyt , the $\varepsilon$-expansion 2022NuPhB.98515990K , and the functional renormalization group Nakayama:2023wrx . As for $\lambda=0$, our result is almost consistent with the other numerical works, while for $\lambda=8$ it shows a clear difference from the one at $\lambda=0$. Thus, it strongly suggests that the dipolar fixed point realized at $\lambda=\infty$ would be in a different universality class from the Heisenberg fixed point.
Figure 1: Comparison plot on the critical exponent $\nu$ (for $\lambda=0$ and $8$) obtained in this paper and the previous studies by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz (colored open symbols), the conformal bootstrap Chester:2020iyt (star-yellow symbol), the $\varepsilon$-expansion (filled symbols), and the functional renormalization group Nakayama:2023wrx (cross-blue-gray symbols). The filled circle-black symbols depict the 1-loop results of the $\varepsilon$-expansion obtained for $\lambda=0$ and $\infty$ in Ref. Aharony_1 and for $\lambda=8$ in this paper whereas the filled square-circle ones are the 3-loop results taken from Ref. 2022NuPhB.98515990K . The two results of the functional renormalization group are obtained with the LPA’ approximation, truncating the potential at the order $n=6$ or $16$ Nakayama:2023wrx .
@@ 4.1 Effectiveness of the local dipolar constraint
We measure the expectation values of the Heisenberg part, $A_{n}:=\sum_{i}\expectationvalue*{\vec{S}_{n}\cdot\vec{S}_{n+\hat{i}}}$, and the local dipolar constraint term, $B_{n}:=\expectationvalue*{(\vec{\nabla}\cdot\vec{S}_{n})^{2}}$, of the Hamiltonian (14), using generated Monte Carlo samples. Here, we take the lattice size $L=40$ and set the temperature to $\beta=0.69019$, $0.35288$, and $0.320374$ for $\lambda=0$, $4$, and $8$, respectively. These values of $\beta$ are indeed the peak positions of the magnetic susceptibility $\chi_{m}$. (See Appendix C for the details of these choices.)
@@
The average values of them over the whole lattice sites are shown in Table 1. We find that the local dipolar constraint term and the Heisenberg term give the same order of contribution to the Hamiltonian given by $\expectationvalue*{H}=\sum_{n}(-A_{n}+\lambda B_{n})$. This means that larger $\lambda$ gives the more strict imposition of the constraint as expected. Furthermore, the results for $\lambda=4$ and $8$ are roughly consistent with the virial theorem.
We further investigate a local property of the dipolar constraint. In the Hamiltonian (14), the dipolar constraint term forces the spin $\vec{S}_{n}$ to satisfy $(\vec{\nabla}\cdot\vec{S}_{n})^{2}=0$ at each site $n$. Here, we measure $B_{n}$ locally at each site $n=(x,0,0)$ as depicted in Figure 2 for $\lambda=0$ (circle-blue symbol), $4$ (square-green), and $8$ (diamond-red).
Figure 2: The results of $B_{n}=\expectationvalue*{(\vec{\nabla}\cdot\vec{S}_{n})^{2}}$ are plotted against the coordinate $x$ on the $x$-axis $n=(x,0,0)$. The lattice size is set to $L=40$. The symbols of circle-blue, square-green, and diamond-red correspond to $\lambda=0$, $4$, and $8$, respectively. The colored horizontal lines depict the average values over the whole lattice sites.
@@ 5.4 Critical exponent: $\nu$
These results of $\nu$ above are compared with the results of previous studies with various methods as shown in Figure 1. Here, we compare the values of $\nu$ for $\lambda=0$, $8$, and $\infty$ obtained by the Monte Carlo Hasenbusch:2020pwj ; Hasenbusch:2011zwv ; Hasenbusch:2000ph ; Ballesteros:1996bd ; Chen:1993zz ; Holm:1993zz ; Peczak:1991zz ; Lau:1989zz , the conformal bootstrap Chester:2020iyt , the $\varepsilon$-expansion Aharony_1 ; 2022NuPhB.98515990K , and the functional renormalization group Nakayama:2023wrx . For reference, the conformal bootstrap predicts $\nu=0.7117(2)$ 444 Here, the error in the bootstrap bound is not one-sigma range but rigorous Chester:2020iyt . and the latest Monte Carlo study Hasenbusch:2020pwj reports $\nu=0.71164(10)$ for the Heisenberg model. Our result at $\lambda=0$ is consistent with them. On the other hand, the result at $\lambda=8$ shows a large deviation from the analytical prediction by the 1-loop $\varepsilon$-expansion. Here, we note that the difference between the 1-loop result of the $\varepsilon$-expansion and the 3-loop result at $\lambda=0$ is larger than the one between the 1-loop one at $\lambda=0$ and $\lambda=8$. It indicates that loop corrections are sizable, at least larger than the $\lambda$-dependence.
@@ 7 Summary and discussion
Our obtained value of the critical exponent $\gamma$ turned out to be smaller than the estimation of the other method. There might exist a theoretical, as well as experimental, interpretation of this apparent smallness (if we assumed that it is indeed different from the dipolar fixed point value $\gamma_{\mathrm{dipolar}}$). [^6] The following interpretation was communicated to us by Slava Rychkov. We would like to thank him for his insight. Recall that our simulation at finite $\lambda$ does not correspond to a true RG fixed point but rather to a crossover point. In such cases, it is argued that the effective susceptibility exponent $\gamma_{\mathrm{eff}}=\frac{d\log\chi}{d\log\beta}$ may show a significant dip away from the fixed point Gimenez-Grau:2023lpz . In our case, the deviation from the RG fixed point is mainly caused by finite $\lambda$, and our definition of $\gamma$ is different from $\gamma_{\mathrm{eff}}$, but it is plausible that the apparent smallness of our obtained $\gamma$ may be attributed to the similar effects. We also note that there is experimental evidence for such an effect in dipolar ferromagnets (see e.g. inbook and reference therein).
@@
Finally, we wish there would be a continuous effort to experimentally verify the dipolar effects in real ferromagnets. [^7] In the Chinese character, the root of magnetism is associated with a stone with benevolence. Why? In ancient Chinese stories, it is said that the way the ferromagnet attracts iron rocks makes it look as if an infant is naturally attracted to his or her mother. Probably the dipolar magnet will embody more benevolence because the magnetism from it is divergenceless, so it appears to be more gentle and embracing. So far in human history, the distinct features of dipolar magnetism have been observed in esoteric materials such as Europium compounds, but the universal nature of the benevolence should also be seen in more conventional materials such as iron or nickel in future experiments, just with more care.
@@ Acknowledgements.
The numerical calculations were carried out on Yukawa-21 at YITP in Kyoto University and the PC clusters at RIKEN iTHEMS. The work of E. I. is supported by JST PRESTO Grant Number JPMJPR2113, JSPS Grant-in-Aid for Transformative Research Areas (A) JP21H05190, JSPS Grant Number JP23H05439, JST Grant Number JPMJPF2221, JPMJCR24I3, and also supported by Program for Promoting Researches on the Supercomputer “Fugaku” (Simulation for basic science: from fundamental laws of particles to creation of nuclei) and (Simulation for basic science: approaching the new quantum era), and Joint Institute for Computational Fundamental Science (JICFuS), Grant Number JPMXP1020230411. The work by YN is in part supported by JSPS KAKENHI Grant Number 21K03581. This work is supported by Center for Gravitational Physics and Quantum Information (CGPQI) at Yukawa Institute for Theoretical Physics.